You want to prepare 50cm³ of 0.1m NaOH from exactly 50%(wt/wt) NaOH solutions.
What volume of the 50% solution is required?
1
Expert's answer
2020-02-20T07:44:09-0500

It is given that 50 % (wt/wt) of NaOH solution its means that 50 g of NaOH is present in 100 g of solution, 1cm3

here it is given that volume of solution = 50 cm3

molarity of solution = .1 M

we know that

Molarity ="\frac{ moles of solute }{volume of solution in (mL)}" "\times" 1000

Molar mass of NaOH = 40 g/mole

here from above information i.e. 50 % wt/wt so we have to convert it into w/v

as mass of water present = 50 g and volume of water = mass/density and we know that

density of water = 1g/cm3

volume of water = 50 ml

as we have to prepare the above molarity solution i.e 50ml

so , concentration in terms of w/v for NaOH = "\frac{50 g of NaOH}{50 cm^3 of solution for dilution} \times 100"

so concetration in terms of w/v is 1 g per 100 ml of solution

so for 1000cm3 of solution we will have 10 g of NaOH

so moles of NaOH will be = "\frac{10}{40} = .25" mole

now use concept

M1V1 =M2 V2

"0.1\times50= V_2\times 0.25"

V2= 20 ml

so, volume of 50 % solution required will be 20 ml